OpenStudy (kellyspeakslouder):

Describe how the graph of y = x2 can be transformed to the graph of the given equation. y = (x+4)2 A. Shift the graph of y = x2 right 4 units. B. Shift the graph of y = x2 up 4 units. C. Shift the graph of y = x2 down 4 units. D. Shift the graph of y = x2 left 4 units.

OpenStudy (kellyspeakslouder):

@welshfella

OpenStudy (kellyspeakslouder):

I need your help! I am far behind in this class and I need to submit two assignments tonight!

OpenStudy (kellyspeakslouder):

@mathstudent55

OpenStudy (kellyspeakslouder):

PLEASE SOMEONE HELP ME

OpenStudy (kellyspeakslouder):

@Nnesha

OpenStudy (kellyspeakslouder):

please anyone

OpenStudy (kellyspeakslouder):

@AloneS please help me

OpenStudy (kellyspeakslouder):

@mathmate

OpenStudy (mathmate):

Any function f(x) may be translated to the \(right\) by h units and up k units by g(x)=f(x-h)+k In your particular case, k=0. For example, to translate f(x)=x^3+x 3 units to the \(left\), we have g(x) =f(x-(-3)) =f(x+3) =(x+3)^3+(x+3) =x^3+3x^2+9x+27 + (x+3) =x^3+3x^2+10x+30 In your case, g(x)=f(x+4), figure out the translation that makes this happen.

OpenStudy (kellyspeakslouder):

B? @mathmate

OpenStudy (kellyspeakslouder):

@mathmate

OpenStudy (kellyspeakslouder):

hello?

OpenStudy (kellyspeakslouder):

@jabez177

OpenStudy (kellyspeakslouder):

@mathmate

OpenStudy (kellyspeakslouder):

@mathmale

OpenStudy (kellyspeakslouder):

@jabez177

OpenStudy (mathmale):

Others might be slightly more inclined to help you if you'd express your exponentiation properly: Correct: "y = x^2 can be transformed to the graph of the given equation. y = (x+4)^2

OpenStudy (kellyspeakslouder):

im correct?

OpenStudy (mathmale):

Rule: If you are given y=f(x) and are asked to graph it, then the function f(x-a) has a graph that is identical to that of f(x), BUT has been shifted "a" units to the right (assuming that "a" is positive. Can you use this fact to answer the original question?

OpenStudy (kellyspeakslouder):

yes

OpenStudy (mathmale):

Can't tell you whether you're correct or not; you haven't posted a response (at least not one that I can see).

OpenStudy (kellyspeakslouder):

I said B lol way up in the comments somewhere

OpenStudy (mathmale):

"y = x^2 can be transformed to the graph of the given equation. y = (x+4)^2" Please draw the graph of y=x^2. Then, using the info I've shared with you, decide how the graph of y = (x+4)^2 will differ from that of y=x^2.

OpenStudy (mathmale):

B is not correct. We'll discuss why, later. For now, please read the Rule I've shared with you, above.

OpenStudy (kellyspeakslouder):

A

OpenStudy (kellyspeakslouder):

@mathmale

OpenStudy (kellyspeakslouder):

@mathmale

OpenStudy (sshayer):

x+4=0 x=-4 so 4 units left.

OpenStudy (kellyspeakslouder):

thank you

OpenStudy (sshayer):

x-5=0 if x=5 , 5 units right

OpenStudy (kellyspeakslouder):

y = x2 + 8

OpenStudy (kellyspeakslouder):

@sshayer what about that one?

OpenStudy (sshayer):

when x=0,y=8 so 8 units up

OpenStudy (sshayer):

if y=x^2-7 when x=0 y=-7 so 7 units down

OpenStudy (kellyspeakslouder):

y = (x+9)2-3

OpenStudy (mathmate):

@kellyspeakslouder More than one helper has shown you the way to figure out the answer. The idea is once you have understood how translation works, you can work out the problem by yourself and be proud about it. Otherwise, you'd need "help" whenever this kind of problems pop up.